# High Gain Fusion Target Design using Generative Artificial Intelligence

### v1

@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; or the Mayer Cluster Expansion [28]. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the finite Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].

### v2

@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; or the Mayer Cluster Expansion [28]. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].
@@ V Control of Collective Systems
The fusion target and drive design still needs to be: optimized with Ubuntu genAI, simulated in 3D with conventional finite-element methods, then experimentally verified. With this being said, this theory already explains ten heretofore unexplained experimental and simulation results.